# Laser Interactions with Gas Jets: EMP Emission and Nozzle Damage

### v1

@@ III.5 Plasma Resistance
where $\nu_{ei}$ is the electron-ion collision frequency. Assuming the electron temperature to be of the order of the ionisation potential, we find $\nu_{ei}\sim 10^{14}$ s-1 and the plasma resistivity is $\eta\simeq 5\times 10^{-6}\,\Omega$ m. This corresponds to a very small resistance of order $0.1\,\Omega$, so as soon as the gas is fully ionised, the plasma potential is applied to the nozzle and a discharge current is produced.
@@ IV EMP emission
Knowing the amplitude of the discharge current, we can also estimate the amplitude of the emitted magnetic field at a distance $r$ as $B\simeq\mu_{0}J_{\text{emp}}/2\pi r$. For a nozzle of height $h_{d}=3$ cm, the peak magnetic field amplitude is $\sim 120$ $\mu$T at a distance of $r=1$ m. The gas discharge is thus expected to produce an electromagnetic pulse with electric field of amplitude $E=cB\approx 36$ kV/m (in the plane wave approximation). This value is higher than the electromagnetic susceptibility threshold for electronics [47].
@@ V Nozzle Damage Mechanism
We identify two mechanisms that could lead to nozzle damage in high-power laser experiments with gas jets: Ohmic heating from a discharge current or bombardment by plasma ions. To evaluate which process is more likely to induce melting, we calculate the energy required to melt a given mass of nozzle material.
@@ V.1 Energy Released through Ohmic Heating
The energy released by Ohmic heating is $J^{2}Rt_{p}$, where $J=Q/t_{p}$ is the discharge current. The resistance $R$ is calculated from the metal resistivity $\eta_{d}$ and the current-carrying surface $A_{d}$ and nozzle length $h_{d}$. Since the discharge is short, the current is confined within a skin depth, $\delta=\sqrt{\eta_{d}/\nu_{\rm emp}\mu_{0}}$. For Cu resistivity $\eta_{d}=1.7\times 10^{-7}\,\Omega$ m and the frequency $\nu_{\rm emp}\approx 2.5$ GHz, the skin depth is $\delta\sim 10\leavevmode\nobreak\ \mu$m. Assuming the current flows along $h_{d}=3$ cm nozzle with a radius of $r_{d}=5$ mm, the current-carrying cross-sectional area is $A_{d}\sim 3\times 10^{-3}$ cm2 and the resistance is $R=\eta_{d}h_{d}/A_{d}\approx 0.01\,\Omega$. The energy released by Ohmic heating with a 4 kA current over $t_{p}\sim 60$ ps is therefore $\approx 10\,\mu$J.
Will the conducting layer melt if it is supplied with $10\,\mu$J of heat energy? The heat conductance length over the heating time is smaller than the skin depth, so the heating is confined within the skin depth. The volume of heated material is therefore $h_{d}A_{d}\sim 10^{-2}$ cm3, which gives a heated mass of $\sim 80$ mg. The available electrostatic energy is not sufficient to heat such a large mass, so it is unlikely that the melting of gas nozzles is caused by Ohmic heating from a discharge current.
@@ VI.1 Ion Acceleration Experiment on VEGA-3
To measure the electron density in the plasma channel, a folding-wave interferometer was used to take optical interferograms of the ionised gas at different temporal delays relative to the drive laser [53]. Fig. 4b shows a He density distribution from the experiment taken $\simeq 150$ ps after the arrival of the laser, revealing a laser channel with average radius $\simeq 70\leavevmode\nobreak\ \mu$m. Modelling the expansion as a Sedov-Taylor cylindrical blast wave [54, 55], $R(t)=\xi_{0}(\mathcal{E}_{b}t^{2}/\rho_{\rm gas}l_{p})^{1/4}$ gives an energy of $\mathcal{E}_{b}\simeq 1.3$ J (taking $\xi_{0}=0.9$ [55]) deposited in the channel over its length of $l_{p}\simeq 0.5$ mm, which is less than 20% the deposited laser energy of 7.3 J. Such a significant difference is explained by the fact that the expanding plasma is essentially collisionless. Plasma ions propagate through the ambient gas and ionise atoms without creating a density compression - only a small portion of the slow ion population contribute to shock wave formation.
2D PIC simulations and experimental interferograms of the gas at the point of laser arrival suggest that, by the time the laser has travelled through the low-density wings of the gas density profile and arrived in the central high-density region, the laser has become significantly defocused. For a gas density of 4.5 bar and a plasma channel radius of $20\leavevmode\nobreak\ \mu$m, the plasma volume is $\sim 6\times 10^{-7}$ cm3, the laser energy deposited in the gas is 7.3 J, the bulk electron temperature is $T_{e}\simeq 77$ keV and the pressure is 60 Mbar (see Tab. 1).
The stopping power of 77 keV electrons in He is $\sim 10$ MeV cm2/g [40], which corresponds to a mean free path $\sim 30$ cm for our gas density. The Debye radius, however, is of order $0.1\leavevmode\nobreak\ \mu$m, so bulk electrons remain in the plasma due to the Coulomb attraction. Plasma expansion is therefore quasi-neutral and Sedov’s self-similar solution for the shock wave applies. Bremsstrahlung losses are negligible on this timescale, so the plasma expands adiabatically.
Assuming that all bulk electron energy is transferred to ions, the average energy of helium ions is about 2320 keV, their velocity is $3\,\mu$m/ps and the expansion time is 150 ps. In this experiment, about 10% of the absorbed laser energy is transferred to hot electrons with an effective temperature of $\sim$1.1 MeV given by the ponderomotive scaling. Then using the plasma charging model described in Sec. II, we estimate a plasma capacitance of 0.0086 pF, a plasma potential of 4.4 MV and 42 nC charge of escaped electrons. The plasma electrostatic energy is about 0.1 J and the escaped electrons have carried away 0.15 J.
@@
Progress of gas ionisation between the laser and nozzle was measured with a $2\omega$ pick-off beam directed transversely to the laser channel for interferometric imaging. Fig. 5 shows interferograms taken at two different times after the arrival of the laser pulse. The gas is N2. Figure 5(a) corresponds to the earliest probing time measured after the laser pulse arrival at $t=0$ ps. The shadow of the gas is visible in both images, suggesting that the entire gas volume can be ionised within 40 ps. The speed of the ionisation front therefore exceeds $\sim 20\,\mu$m/ps, which is consistent with gas ionisation by electrical breakdown.
@@ VI.2 EMP Experiment on VEGA-3
A second experiment was conducted on VEGA-3 with a view to characterising the EMP and ions produced in laser-gas interactions. The laser energy before compressor was $\sim 30$ J and the pulse duration was 30 fs. Based on images of the focal spot taken at low energy, 21% of the energy on-target was contained within the first Airy disk[50] for an on-target intensity of order $\sim 10^{20}$ W cm-2 at best compression, similar to the intensity reported for the experiment in the previous section. The laser contrast was $10^{-12}$ with no pre-pulses. A Prodyn RB-230(R) radiation-hardened B-dot probe was used to measure the amplitude of the magnetic field at different distances from the gas jet, while two diamond time of flight detectors and a Thomson parabola were used to monitor charged particle emission. Signal degradation due to EMP pick-up in the unshielded oscilloscope prevented the easy identification of a photo-peak in the time of flight data, so quantitative analysis of the ion spectrum is not possible. B-dot signals were recorded on a Rohde&Schwarz RTO64 digital oscilloscope with 2 GHz bandwidth and 10 GS/s sample rate. The B-dot probe was oriented so that its axis was parallel to the ground and sensitive to an azimuthal magnetic field relative to the nozzle axis.
@@ VI.2.1 Comparison with EMP Emission Model
The magnitude of the magnetic field is of order $5-10\,\mu$T at a distance of $\sim 2$ m from the jet. Significant shot-to-shot variations are due to adjustments in the laser focal position inside the gas to optimise the laser coupling to the target. There are insufficient data points to identify an unambiguous scaling with distance, but they can allow us to estimate the charge accumulated in the gas by the laser ejection of hot electrons if we assume a dipolar radiation field. The solid line represents a least squares fit to the data points with Eq. (1) from Ref. [46], assuming an antenna height of $h=3$ cm (consistent with the vertical height of the jet nozzle in Fig. 6) and leaving the target charge $Q$ as a free parameter. The best-fit charge is $\sim 20$ nC in this case. This is in agreement with the model presented in Sec. II, which predicts a charge of $\sim 40$ nC for the VEGA-3 experimental conditions.
Peaks in the EMP Fourier spectrum constrain the EMP emission mechanism. The average of the EMP Fourier spectra for shots with the gas jet targets reveals multiple prominent resonances between $\nu_{\rm emp}\sim 1.2$ GHz and 1.9 GHz (not shown here). Assuming antenna emission from a monopole of height $c/4\nu_{\rm emp}$, the strongest resonances correspond to heights of $\sim 4-5$ cm - similar to the dimensions of the nozzle. Note that our measurements were only sensitive to emissions from 0.4 GHz to 2 GHz, and there were other large metallic objects in the chamber that might contribute to the EMP.
@@ VII Conclusions
Further dedicated experiments are needed to confirm the model presented in this paper, in particular by measuring the emitted ion and electron spectra over a large solid angle. Measuring the average ion energy will allow us to estimate the discharge time and plasma expansion velocity. It is also important to measure the laser-ejected charge since 10 times more hot electrons imply 10 times stronger EMP fields. Direct measurement of a discharge current in the nozzle and characterisation of the EMP field structure is essential to confirm that the EMP is related to antenna emission from the nozzle rather than large-scale plasma oscillations or chamber resonant modes [6, 60].
Experimental validation of the nozzle damage mechanism requires measurement of the ion energy distribution in the radial direction, in the energy range below 1 MeV. This might also be combined with a study of laser damage as the laser is focused at various different heights above the gas nozzle. Switching from metallic to ceramic nozzles would be useful for assessing the impact of nozzle material.
@@ Acknowledgements.
The authors would like to thank the staff at the VEGA-3 Laser Facility for their hard work over the course of several experiments. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No. 101052200 — EUROfusion). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. We acknowledge GENCI for providing us access to the Joliot-Curie supercomputer (grants 2021-A0130512993 and 2022-A0130512993). We also acknowledge Grant PID2021-125389OA-I00, funded by MCIN/ AEI / 10.13039/501100011033 / FEDER, UE and “ERDF A way of making Europe”, funded by the European Union. V. Ospina-Bohórquez and C. Vlachos acknowledge support from the LIGHT S&T Graduate Program (PIA3 Investment for the Future Program, ANR-17-EURE-0027). This work received funding from the European Union’s Horizon 2020 research and innovation program through the European IMPULSE project under grant agreement No 871161 and from LASERLAB-EUROPE V under grant agreement No 871124. It is published as part of the international project “PMW”, co-financed by the Polish Ministry of Science and Higher Education within the framework of the scientific financial resources for 2021-2022 under the contract no 5205/CELIA/2021/0 (project CNRS no 239915).

### v2

@@ III.5 Plasma Resistance
where $\nu_{ei}$ is the electron-ion collision frequency. Assuming the electron temperature to be of the order of the ionisation potential, we find $\nu_{ei}\sim 10^{14}$ s-1 and the plasma resistivity is $\eta\simeq 5\times 10^{-6}\,\Omega$ m. This corresponds to a very small resistance of order $0.1\,\Omega$. A path to ground is established when the gas is fully ionised; the potential drop across the nozzle is of comparable magnitude to the plasma potential and a discharge current is produced. The accumulated plasma charge and discharge time puts a limit on the maximum discharge current of order $\sim 10$ kA, as discussed in the next section.
@@ IV EMP emission
Knowing the amplitude of the discharge current, we can also estimate the amplitude of the emitted magnetic field at a distance $r$ as $B\simeq\mu_{0}J_{\text{emp}}/2\pi r$. This is equivalent to the maximum B-field (measured over all possible emission angles) in the far-field of a dipole antenna[51]. For a nozzle of height $h_{d}=3$ cm, the peak magnetic field amplitude is $\sim 120$ $\mu$T at a distance of $r=1$ m. The gas discharge is thus expected to produce an electromagnetic pulse with electric field of amplitude $E=cB\approx 36$ kV/m (in the plane wave approximation). This value is higher than the electromagnetic susceptibility threshold for electronics [52].
@@ V Nozzle Damage Mechanism
We identify two mechanisms [^1] UV, soft x-rays and thermal electrons can also couple to the nozzle and produce heating. Hard x-rays will deposit their energy volumetrically and are less important. The energy in UV and x-rays is relatively small, however (see Sec. III.1), while the thermal electrons travel with the ions and have a lower stopping power[42], producing heating in conjunction with the ions. that could lead to nozzle damage in high-power laser experiments with gas jets: Ohmic heating from a discharge current or bombardment by plasma ions. To evaluate which process is more likely to induce melting, we calculate the energy required to melt a given mass of nozzle material.
@@ V.1 Energy Released through Ohmic Heating
The energy released by Ohmic heating is $J^{2}Rt_{p}$, where $J=Q/t_{p}$ is the discharge current. The resistance $R$ is calculated from the metal resistivity $\eta_{d}$ and the current-carrying surface $A_{d}$ and nozzle length $h_{d}$. Since the discharge is short, the current is confined within a skin depth, $\delta=\sqrt{\eta_{d}/(\nu_{\rm emp}\mu_{0})}$. For Cu resistivity $\eta_{d}=1.7\times 10^{-7}\,\Omega$ m and the frequency $\nu_{\rm emp}\approx 2.5$ GHz, the skin depth is $\delta\sim 10\leavevmode\nobreak\ \mu$m. Assuming the current flows along $h_{d}=3$ cm nozzle with a radius of $r_{d}=5$ mm, the current-carrying cross-sectional area is $A_{d}\sim 3\times 10^{-3}$ cm2 and the resistance is $R=\eta_{d}h_{d}/A_{d}\approx 0.01\,\Omega$. The energy released by Ohmic heating with a kA-level current over $t_{p}\sim 2h/c\approx 3$ ps is therefore $\lesssim 0.1\,\mu$J.
Will the conducting layer melt if it is supplied with $0.1\,\mu$J of heat energy? The volume of heated material is $h_{d}A_{d}\sim 10^{-2}$ cm3, which gives a heated mass of $\sim 80$ mg. The available electrostatic energy is not sufficient to heat such a large mass, so it is unlikely that the melting of gas nozzles is caused by Ohmic heating from a discharge current.
@@ VI.1 Ion Acceleration Experiment on VEGA-3
To measure the electron density in the plasma channel, a folding-wave interferometer was used to take optical interferograms of the ionised gas at different temporal delays relative to the drive laser [57]. Figure 4b shows a He density distribution from the experiment taken $\simeq 150$ ps after the arrival of the laser, revealing a laser channel with average radius $\simeq 70\leavevmode\nobreak\ \mu$m. Modelling the expansion as a Sedov-Taylor cylindrical blast wave [58, 59], $R(t)=\xi_{0}(\mathcal{E}_{b}t^{2}/\rho_{\rm gas}l_{p})^{1/4}$ gives an energy of $\mathcal{E}_{b}\simeq 1.3$ J (taking $\xi_{0}=0.9$ [59]) deposited in the channel over its length of $l_{p}\simeq 0.5$ mm, which is less than 20% the deposited laser energy of 7.3 J. Such a significant difference is explained by the fact that the expanding plasma is essentially collisionless. Plasma ions propagate through the ambient gas and ionise atoms without creating a density compression - only a small portion of the slow ion population contribute to shock wave formation.
2D PIC simulations in CALDER[25, 29] and experimental interferograms of the gas at the point of laser arrival suggest that, by the time the laser has travelled through the low-density wings of the gas density profile and arrived in the central high-density region, the laser has become significantly defocused. For a gas density of 4.5 bar and a plasma channel radius of $20\leavevmode\nobreak\ \mu$m, the plasma volume is $\sim 6\times 10^{-7}$ cm3, the laser energy deposited in the gas is 7.3 J, the bulk electron temperature is $T_{e}\simeq 140$ keV and the pressure is $\sim$50 Mbar (see Tab. 1).
The stopping power of 140 keV electrons in He is $\sim 3$ MeV cm2/g [42], which corresponds to a mean free path $>50$ cm for our gas density. The Debye radius, however, is of order $0.1\leavevmode\nobreak\ \mu$m, so bulk electrons remain in the plasma due to the Coulomb attraction and the plasma expansion is quasi-neutral. Bremsstrahlung losses are negligible on this timescale, so the plasma expands adiabatically.
Assuming that all bulk electron energy is transferred to ions, the average energy of helium ions is about 0.43 MeV, their velocity is $\sim 5\,\mu$m/ps and the expansion time is 100 ps. In this experiment, about 10% of the absorbed laser energy is transferred to hot electrons with an effective temperature of $\sim$1.1 MeV given by the ponderomotive scaling. Then using the plasma charging model described in Sec. II, we estimate a plasma capacitance of 0.0086 pF, a plasma potential of 4.4 MV and 42 nC charge of escaped electrons. The plasma electrostatic energy is about 0.1 J and the escaped electrons have carried away $\sim$0.15 J.
@@
Progress of gas ionisation between the laser and nozzle was measured with a $2\omega$ pick-off beam directed transversely to the laser channel for interferometric imaging. Figure 5 shows interferograms taken at two different times after the arrival of the laser pulse. The gas is N2, here. Figure 5(a) corresponds to the earliest probing time measured after the laser pulse arrival at $t=0$ ps. The shadow of the gas is visible in both images, suggesting that the entire gas volume can be ionised within 40 ps. The speed of the ionisation front therefore exceeds $\sim 20\,\mu$m/ps, which is consistent with gas ionisation by electrical breakdown.
@@ VI.2 EMP Experiment on VEGA-3
A second experiment was conducted on VEGA-3 with a view to characterising the EMP and ions produced in laser-gas interactions. The laser energy before compressor was $\sim 30$ J and the pulse duration was 30 fs. Based on images of the focal spot taken at low energy, 21% of the energy on-target was contained within the first Airy disk[23] for an on-target intensity of order $\sim 10^{20}$ W cm-2 at best compression, similar to the intensity reported for the experiment in the previous section. The laser contrast was $10^{-12}$ at $0.1$ ns, with no significant pre-pulses[23]. A Prodyn RB-230(R) radiation-hardened B-dot probe was used to measure the amplitude of the magnetic field at different distances from the gas jet, while two diamond time of flight detectors and a Thomson parabola were used to monitor charged particle emission. Signal degradation due to EMP pick-up in the unshielded oscilloscope prevented the easy identification of a photo-peak in the time of flight data, so quantitative analysis of the ion spectrum is not possible. B-dot signals were recorded on a Rohde&Schwarz RTO64 digital oscilloscope with 2 GHz bandwidth and 10 GS/s sample rate. The B-dot probe was oriented so that its axis was parallel to the ground and sensitive to an azimuthal magnetic field relative to the nozzle axis.
@@ VI.2.1 Comparison with EMP Emission Model
The magnitude of the magnetic field is of order $5-10\,\mu$T at a distance of $\sim 2$ m from the jet. Significant shot-to-shot variations are due to adjustments in the laser focal position inside the gas to optimise the laser coupling to the target. There are insufficient data points to identify an unambiguous scaling with distance, but they can allow us to estimate the charge accumulated in the gas by the laser ejection of hot electrons if we assume a dipolar radiation field. The solid line represents a least squares fit to the data points with Eq. (1) from Ref. [51], assuming an antenna height of $h_{d}=3$ cm (consistent with the vertical height of the jet nozzle in Fig. 6) and leaving the target charge $Q$ as a free parameter. The best-fit charge is $\sim 20$ nC in this case. This is in agreement with the model presented in Sec. II, which predicts a charge of $\sim 40$ nC for the VEGA-3 experimental conditions.
Peaks in the EMP Fourier spectrum constrain the EMP emission mechanism. The average of the EMP Fourier spectra for shots with the gas jet targets reveals multiple prominent resonances between $\nu_{\rm emp}\sim 1.2$ GHz and 1.9 GHz (not shown here). Assuming antenna emission from a monopole of height $c/4\nu_{\rm emp}$, the strongest resonances correspond to heights of $\sim 4-5$ cm - similar to the dimensions of the nozzle. Note that our measurements were only sensitive to emissions from 0.4 GHz to 2 GHz and there were other large metallic objects in the chamber that might contribute to the EMP.
@@ VII Conclusions
Further dedicated experiments are needed to confirm the model presented in this paper, in particular by measuring the emitted ion and electron spectra over a large solid angle. Measuring the average ion energy will allow us to estimate the discharge time and plasma expansion velocity. It is also important to measure the laser-ejected charge since 10 times more hot electrons imply 10 times stronger EMP fields. Direct measurement of a discharge current in the nozzle and characterisation of the EMP spectrum is essential to confirm that the EMP is related to antenna emission from the nozzle.
EMP sources couple with the chamber and objects within it, exciting resonant modes[6]. These modes are excited differently according to the nature of the driver and may last much longer than the source itself, depending on their relaxation time. We have discussed a plasma discharge as one important source of EMP, but the electron beam ejected from the plasma can also induce currents and secondary electromagnetic fields in the nozzle and surrounding chamber [64]. The intensity of the induced current is much smaller than that of the escaped electrons, however, and its impact on the nozzle damage is expected to be small.
Experimental validation of the nozzle damage mechanism requires measurement of the ion energy distribution in the radial direction, in the energy range below 1 MeV. This might also be combined with a study of nozzle damage as the laser is focused at various different heights above the gas nozzle. Switching from metallic to ceramic nozzles would be useful for assessing the impact of nozzle material.
@@ Acknowledgements.
The authors would like to thank the staff at the VEGA-3 Laser Facility for their hard work over the course of several experiments. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No. 101052200 — EUROfusion). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. We acknowledge GENCI for providing us access to the Joliot-Curie supercomputer (grants 2021-A0130512993 and 2022-A0130512993). We also acknowledge Grant PID2021-125389OA-I00, funded by MCIN/ AEI / 10.13039/501100011033 / FEDER, UE and “ERDF A way of making Europe”, funded by the European Union. V. Ospina-Bohórquez and C. Vlachos acknowledge support from the LIGHT S&T Graduate Program (PIA3 Investment for the Future Program, ANR-17-EURE-0027). This work received funding from the European Union’s Horizon 2020 research and innovation program through the European IMPULSE project under grant agreement No 871161 and from LASERLAB-EUROPE V under grant agreement No 871124. It is published as part of the international project “PMW”, co-financed by the Polish Ministry of Science and Higher Education within the framework of the scientific financial resources for 2021-2022 under the contract no 5205/CELIA/2021/0 (project CNRS no 239915). Finally, we acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program “GPR LIGHT”.